OpenStudy (anonymous):

If G(x) = (x + 4)2, and F(x) = x2 + 8x, what is the function G(F(x))?

OpenStudy (anonymous):

you need to write it more clearly, is x2 =x^2, or x*2?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

ok,

OpenStudy (anonymous):

x^4+16x^3+74X^2+64x+16

jimthompson5910 (jim_thompson5910):

\[G(x) = (x+4)^2\] \[G(F(x)) = (F(x)+4)^2\] \[G(F(x)) = (x^2+8x+4)^2\] \[G(F(x)) = (x^2+8x+4)(x^2+8x+4)\] \[G(F(x)) = x^2(x^2+8x+4)+8x(x^2+8x+4)+4(x^2+8x+4)\] \[G(F(x)) = x^4+8x^3+4x^2+8x^3+64x^2+32x+4x^2+32x+16\] \[G(F(x)) = x^4+16x^3+72x^2+64x+16\]

OpenStudy (anonymous):

yep, thats right, i made a small mistake in 74x^2, its 72x^2